Lecture #2: Foundations of Probability
MATH 321 – Mathematical Statistics I
Torin Quinlivan
Knox College
October 17, 2026
What is Probability?
A probability \(p\) of an event is a number \(0 \le p \le 1\).
The frequentist interpretation is that \(p\) is the proportion of times the event will occur if I repeat the experiment over and over.
The Bayesian interpretation is that \(p\) is my degree of belief that the event will occur, given what I believe to be true. (We can update this belief as we get more information.)
What is Probability?
Toss a coin. What is the probability of getting a heads?
Throw a 6-sided die. What is the probability of getting a “1”?
Draw five cards from a deck. What is the probability of drawing two aces?
Careful with the wording! Does this question mean:
What is the probability of drawing exactly two aces?
What is the probability of drawing at least two aces?
Complement Rule
Corollary (Complement Rule):
\[
\forall A \in \mathcal{P}(\Omega) \; (P(\overline{A}) = 1 - P(A)).
\]
What does this mean in English?
For all events \(A\), the probability of the complement of the event, \(\overline{A}\), is \(1\) minus the probability of the event \(A\).
This makes intuitive sense. If the Bears have a \(46\%\) chance of winning on Sunday, they have a \(100\% - 46\% = 54\%\) chance of not winning.
Inclusion/Exclusion Principle
Corollary (Inclusion/Exclusion Principle):
\[
\forall A, B \in \mathcal{P}(\Omega) \; (P(A \cup B) = P(A) + P(B) - P(A \cap B)).
\]
What does this mean in English?
For any two events \(A\) and \(B\), the probability of the union of the two events is equal to the probability of \(A\) plus the probability of \(B\) minus the probability of the intersection of \(A\) and \(B\) (because it’s being double counted).
Inclusion/Exclusion Principle
Corollary (Inclusion/Exclusion Principle):
\[
\forall A, B \in \mathcal{P}(\Omega) \; (P(A \cup B) = P(A) + P(B) - P(A \cap B)).
\]
This can be extended to more than two events, with the pluses and minuses alternating:
\[
\begin{aligned}
P \left( \bigcup A_i \right) &= P(A_1) +P(A_2) + \dots + P(A_n) \\
&\quad - \, P(A_1 \cap A_2) - P(A_1 \cap A_3) - \dots - P(A_{n-1} \cap A_n) \\
&\quad + \, P(A_1 \cap A_2 \cap A_3) + P(A_1 \cap A_2 \cap A_4) + \dots \\
&\quad + \, P(A_{n-2} \cap A_{n-1} \cap A_n) - P(A_1 \cap A_2 \cap A_3 \cap A_4) \\
&\qquad \vdots \\
\end{aligned}
\]
Examples
Example: Consider rolling a fair six-sided die.
In our formal mathematical notation, let \(\Omega = \{ 1, 2, 3, 4, 5, 6 \}\) and for \(x \in \Omega\), \(P(x) = \dfrac{1}{6}\).
What is the probability of rolling a \(1\)?
What is the probability of not rolling a \(1\)?
What is the probability of rolling an even number?
What is the probability of rolling an odd number?
Equally Likely Outcomes
If \(\lvert \Omega \rvert\) is finite and all atomic outcomes in \(\Omega\) are equally likely, then for \(x \in \Omega\) and \(A \subseteq \Omega\): \[
P(x) = \frac{1}{\lvert \Omega \rvert} \qquad \text{ and } \qquad P(A) = \frac{\lvert A \rvert}{\lvert \Omega \rvert}.
\]
The probability of \(A\) is the number of outcomes in \(A\) divided by the total number of possible outcomes.
A uniform discrete probability space has \(N\) outcomes, each of which occurs with probability \(\frac{1}{N}\).
Lots of probability spaces are uniform and you can evaluate their probabilities using the probability axioms and counting formulas.
Equally Likely Outcomes Examples
Example: Consider a standard deck of 52 cards and suppose a hand of five cards is drawn.
What is the probability of drawing a hand with two aces?
What is the probability of drawing a pair of aces and three kings?
What is the probability of drawing a full house?
Probability Distributions
Random Variables
A random variable \(X\) on \(\Omega\) is a function \(X : \Omega \to \mathbb{R}\). Since each \(x \in \Omega\) has a probability associated with it, each \(X(x)\) has a probability as well: \(P(X = a) = P(\{ x \in \Omega : X(x) = a \})\).
Probability Mass Function
The probability mass function of a discrete random variable gives the probability that it takes on each value in its range: \(f(x) = P(X=x)\).
Probability Distribution Example
Suppose we roll two dice and take the sum of the values. The probability distribution for \(X\) is:
| \(\boldsymbol{x}\) |
2 |
3 |
4 |
5 |
6 |
7 |
8 |
9 |
10 |
11 |
12 |
| \(\boldsymbol{P(x)}\) |
\(\dfrac{1}{36}\) |
\(\dfrac{2}{36}\) |
\(\dfrac{3}{36}\) |
\(\dfrac{4}{36}\) |
\(\dfrac{5}{36}\) |
\(\dfrac{6}{36}\) |
\(\dfrac{5}{36}\) |
\(\dfrac{4}{36}\) |
\(\dfrac{3}{36}\) |
\(\dfrac{2}{36}\) |
\(\dfrac{1}{36}\) |
What is the probability the sum is greater than 9?
What is the probability the sum is odd?
Probability Distributions
Distribution functions give a way to compute probability when outcomes aren’t equally likely.
Distribution functions come in two flavors: probability mass (or density) functions and cumulative distribution functions, which represents \(P(X \le x)\).
Cumulative distribution functions are most often used when dealing with continuous random variables (e.g., the normal distribution functions) because probabilities can be evaluated with integrals.
Probability distributions give us a way to classify families of similar random variables. The study of families of probability distributions occupies a large part of probability theory (and statistics).
The Binomial Distribution
Suppose an experiment that has two possible outcomes (success or failure); is (independently) repeated \(n\) times (trials); and that the probability of success (each time) is \(p\). Let \(X\) be a random variable which counts the number of successes that occur among these \(n\) trials. Then the probability mass function is:
\[
f(x) = P(X = x) = \binom{n}{x} p^x (1-p)^{n-x}, \quad x = 0, 1, 2, \dots, n
\]
Example: Suppose a loaded coin which has probability of heads \(p = .7\) is tossed 10 times.
What is the probability of getting exactly \(8\) heads?
What is the probability of getting at least \(8\) heads?